OpenStudy (anonymous):

any help with this one. find the exact value of the expression cos(-120deg)

OpenStudy (unklerhaukus):

\[\cos(-\theta)=\cos(\theta)\]

OpenStudy (anonymous):

cos(-120)=cos(240) =  \[\sqrt{3}/2\]